OpenStudy (anonymous):

linear combination questions : 2x-y=5 2x+y=3 x+y=6 3x - 3y= 18

OpenStudy (texaschic101):

2x - y = 5 2x + y = 3 -----------add 4x = 8 x = 2 sub 2 in for x in either equation to find y 2x - y = 5 2(2) - y = 5 4 - y = 5 4 - 5 = y -1 = y solution (2,-1) ================== 3x - 3y = 18 reduces to x - y = 6 now we have : x - y = 6 x + y = 6 -----------add 2x = 12 x = 6 now sub 6 in for x in either equation to find y x + y = 6 6 + y = 6 y = 6 - 6 y = 0 solution is (6,0) =================== if you have any questions, please ask :)